# Configurable DOA Estimation using Incremental Learning

### v2

@@ III-C Incremental learning setting
  • Finetune: Starting with a model pre-trained on the initial task, we sequentially fine-tune it on each subsequent task. This approach adapts the model incrementally, but may be susceptible to forget previously learned information when moving to new tasks each time.
@@
  • Joint: A single model is trained on the entire dataset, encompassing all microphone spacings together. This method aims to learn a general representation that performs well across all the tasks.
@@
  • Ensemble: This involves training five separate models, each on a different subset of the data, with each model specializing in a particular microphone spacing.
III-D Implementation details
@@
For training each task within our DOA-PNN framework, we consistently apply 100 epochs with an early stopping criterion set to activate after 20 epochs if there’s no improvement in performance. We employ the AdamW [28] optimizer, a variant of the Adam optimizer that includes weight decay for regularization, with a learning rate of 0.001. Additionally, we apply a learning rate scheduler, specifically StepLR, to adjust the learning rate during training.
For the STFT computations, we set the frame size to 32 milliseconds and the hop size to 10 milliseconds. The frequency range for the STFT is confined between a minimum of 100 Hz and a maximum of 8000 Hz, using an n_fft value of 512, which determines the length of the windowed signal after padding with zeros. This STFT configuration is designed to capture a wide range of audio frequencies suitable for accurately determining the DOA. Gaussian function constant $\sigma$ we set as 8 in our experiments.
TABLE II: Performance comparison across different DOA estimation methods, highlighting MAE, ACC within various degree tolerances, and model parameters.
Methods MAE ACC(±$\pm 5$) ACC(±$\pm 10$) ACC(±$\pm 15$) Params
Finetune 17.8 33.9 48.0 58.4 0.9M
Joint 11.9 54.0 66.7 74.3 0.9M
Ensemble 7.6 78.6 84.8 87.5 4.8M
Proposed DOA-PNN 10.1 72.9 82.7 86.1 1.6M
IV Results and Discussions
@@
We test the methods for DOA estimation across varying microphone spacings as Fig. 2, we can observe how each method adapts to the acoustic challenge. The finetune method incrementally trained from 5cm to 9cm only performs well for the recent task 9 cm indicating a forgetting across previously learned configurations. On looking to the performance for joint training, we find that it is more resilience compared to finetune, yet it does not maintain high performance as spacing widens. It is possible that the single-model approach is not able to capture the nuances needed for localization at variable spacings. The ensemble method demonstrates a robust performance across different spacings as the upper bound. It benefits from training distinct models for each spacing scenario.
Our proposed DOA-PNN shows a promising ability to sustain higher accuracy levels, particularly within the most challenging ±$\pm 5$-degree margin. It implies that our progressive learning strategy, which builds upon knowledge from previous configurations transfers well to new spacing scenarios. The proposed DOA-PNN offers a balance between accuracy and model complexity, suggesting it to be an efficient choice for real-world applications where microphone spacing can vary.
On examining the comparative results of different methods for DOA estimation as given in Table II, we find each method has its strengths and limitations. The finetune approach, while the most straightforward in terms of adapting pre-existing models to new tasks, seems to lag in precision. It presents the highest MAE, suggesting that it might not preserve the nuances of DOA estimation in different microphone spacing as effectively as the other methods. The joint training method improves upon finetune significantly, indicating that learning from a collective dataset of all tasks offers a better understanding of the DOA estimation problem. However, this method still does not reach the top performance in accuracy, especially in ±$\pm 5$-degree range. Ensemble, which employs a separate model for each condition, stands out in accuracy. It’s clear that specializing in individual conditions allows for a finer understanding of each specific DOA estimation task, but this comes at the cost of a larger overall model size, as reflected by the parameter count.
Our Proposed DOA-PNN method maintains a balance between precision and efficiency. It achieves comparable accuracy, nearly matching the best-performing method in the ±$\pm 15$-degree range and outperforming it within the ±$\pm 10$-degree margin. It achieves this while maintaining a moderate model size. This indicates that our progressive approach, which incorporates learning from successive tasks, effectively consolidates knowledge without increasing complexity.
@@
We foresee several avenues for future work. While we currently focus on two-microphone linear arrays to estimate the azimuth angle, we plan to expand our research to encompass more complex acoustic configurations and to address elevation angle estimation for a full spatial representation [29]. Moreover, integrating noise robustness factors into DOA estimation is a critical next step, acknowledging that real-world applications often operate in acoustically challenging environments [30]. These enhancements will aim to bolster the practicality of DOA-PNN, providing comprehensive solutions for 3D sound localization in noisy, real-life settings.
V Conclusions
@@
In this study, we presented DOA-PNN, an approach that advances the direction of arrival estimation within continual learning. DOA-PNN synergizes task-specific network instantiation with progressive learning principles, addressing the diversity in microphone configurations with a novel scaling mechanism and lateral connections for knowledge retention. Demonstrated across simulated acoustic settings, DOA-PNN consistently achieved high accuracy, underscoring its capacity for real-world applicability in dynamic auditory scenes. The model’s efficiency and adaptability suggest its potential as a foundational framework for future research in sound localization and acoustic signal processing, encouraging the development of sophisticated, scalable audio analysis technologies.

### v3

@@ IV-C Incremental learning setting
  • Finetune: Starting with a model pre-trained on the initial task, we sequentially fine-tune it on each subsequent task. This approach adapts the model incrementally but may be susceptible to forgetting previously learned information when moving to new tasks each time.
@@
  • Joint: A single model is trained on the entire dataset, encompassing all conditions. This method aims to learn a general representation that performs well across all tasks. However, it can not handle large domain mismatches.
@@
  • Multi-Condition (MC): This involves training separate models for each condition on a different subset of the data, with each model specializing in a particular task. In incremental task setting, the performance can never surpass the MC as it already knows the whole target classes the first time and ensemble multiple models.
  • Elastic Weight Consolidation (EWC) [35]: is a continual learning method that uses a quadratic penalty to restrict changes to parameters critical for previous tasks. These critical parameters are identified using an approximation based on the Fisher Information Matrix.
  • Learning without Forgetting (LwF) [36]: is a method that ensures previously learned knowledge is retained while training on new tasks. It uses knowledge distillation, where the model’s predictions on old tasks are treated as “teacher” during the new task training.
  • Memory Aware Synapses (MAS) [37]: is similar to EWC but identifies critical parameters by measuring their impact on output predictions.
IV-D Implementation details
@@
In our CL-SSL framework, each task is trained for up to 100 epochs, with early stopping triggered after 20 epochs without performance improvement. We use the AdamW optimizer with weight decay, a learning rate of 0.001, and the StepLR scheduler for adjustments during training. For STFT computations, we use a frame size of 32 ms, a hop size of 10 ms, and a frequency range of 100 Hz to 8000 Hz, with n_fft set to 512. The Gaussian constant $\sigma$ is set to 8 in our experiments.
V Results and Discussions
V-A Results on simulated dataset
@@
We first test the methods for SSL across varying microphone spacing and report the results in Fig. 2 and Table II. For the non-CL methods, we observe the finetune method incrementally trained only performs well for the first task indicating a forgetting across learned tasks. It presents the highest MAE, suggesting that it might not preserve the nuances of SSL in different microphone spacing as effectively as the other methods in Table II. Observing the performance of joint training, we find that it is more resilient compared to finetune, yet it does not maintain a high performance similar to the MC method. It is possible that the single-model approach is unable to capture the nuances needed for localization at variable spacing. The MC method shows consistent performance across different spacings due to the training of separate models for each task. However, it demands five times more memory and training time, making it less practical for real-world applications. As the number of tasks grows, the memory and time requirements increase linearly.
For CL methods, all methods outperform finetuning, demonstrating their effectiveness in reducing forgetting. However, their performance remains inferior to that of joint training and MC. In addition, from Fig. 2, it is observed that the performance of EWC and MAS drops dramatically when the number of learning tasks increases. This gap may result due to inappropriate regularization, which limits the model’s ability to learn new tasks while preserving prior knowledge. Among the reference CL methods, LwF performs better, likely because its use of the model’s predictions of old tasks enables knowledge sharing across tasks.
Our proposed CL-SSL shows a promising ability to obtain higher accuracy levels, particularly within the most challenging ±$\pm 5$-degree margin as shown in Fig. 2. It achieves comparable accuracy, nearly matching the best-performing method in the ±$\pm 15$-degree range in Table II. This indicates that our continual learning strategy effectively transfers knowledge from previous configurations to new spacing scenarios. By employing the layer scaling mechanism, we maintain a parameter count of just 1.6M—lower than LwF—while achieving significantly better performance. This highlights the proposed CL-SSL as a balanced solution between accuracy and model complexity, making it an efficient choice for SSL applications with varying microphone spacings.
TABLE II: Performance comparison in ACC (%) across different methods on the simulated dataset, highlighting MAE. The performance is calculated after all tasks are learned.
Methods MAE ACC(±$\pm 5^{\circ}$) ACC(±$\pm 10^{\circ}$) ACC(±$\pm 15^{\circ}$) Params
Finetune 26.5 12.7 23.8 34.7 0.9M
Joint 13.0 50.2 64.1 72.2 0.9M
MC 7.6 77.6 85.6 88.4 4.5M
EWC [35] 23.5 25.4 38.7 51.4 0.9M
LwF [36] 22.1 35.5 47.4 58.9 1.8M
MAS [37] 23.3 24.8 36.3 48.0 0.9M
Proposed CL-SSL 10.1 71.5 83.0 87.0 1.6M
V-B Results on real-world dataset
As discussed, we further evaluate the proposed CL-SSL method on a noisy real-world dataset. Noise is added based on SNR values randomly selected from [-5 dB, 0 dB, 5 dB, 10 dB, Clean], representing different acoustic environments, with task order randomly shuffled for practicality. The results compared with other baselines are summarized in Table III. We observe better performance across all methods on the LOCATA dataset due to its more stable environment, including consistent room size, microphone center, and room impulse response. The finetune method remains the lower bound among non-CL methods, indicating that catastrophic forgetting negatively impacts performance across varied acoustic environments due to domain mismatches. Joint training achieves relatively better results on real-world data than on the simulated data (Table II), likely because optimizing for diverse acoustic environments is easier than for varying microphone configurations. The MC method, serving as the upper bound, performs well but at the cost of significantly higher memory usage.
Among the CL baselines, EWC, LwF, and MAS show comparable performance and outperform the finetune method, suggesting that these approaches effectively mitigate forgetting while balancing adaptation to new tasks. Our proposed CL-SSL method demonstrates less forgetting compared to other CL methods and performs closely to the MC method. When compared to joint training, often regarded as the upper bound in many applications, CL-SSL achieves comparable results despite not having access to all tasks in advance. This suggests that our method has greater potential to perform well in unseen acoustic environments. Notably, in the most challenging ±$\pm 5$-degree scenario, CL-SSL even surpasses joint training. This advantage may stem from its ability to adaptively scale parameters and focus on critical features, enabling better fine-grained localization under strict error tolerances.
TABLE III: Performance comparison in ACC (%) across different methods on the real-world dataset, highlighting MAE. The performance is calculated after all tasks are learned.
Methods MAE ACC(±$\pm 5^{\circ}$) ACC(±$\pm 10^{\circ}$) ACC(±$\pm 15^{\circ}$) Params
Finetune 17.2 43.0 58.6 67.7 0.9M
Joint 5.2 68.4 78.9 86.8 0.9M
MC 4.3 78.9 89.4 94.7 4.5M
EWC [35] 9.3 47.3 68.4 81.5 0.9M
LwF [36] 8.9 50.0 65.8 81.8 1.8M
MAS [37] 10.8 48.8 67.1 77.4 0.9M
Proposed CL-SSL 5.1 72.9 78.9 86.5 1.6M
V-C Future Directions
@@
We foresee several avenues for future work. While we currently focus on two-microphone linear arrays to estimate the azimuth angle, we plan to expand our research to encompass more complex acoustic configurations [38] and to address elevation angle estimation for a full spatial representation.
VI Conclusions
@@
In this study, we presented CL-SSL, an approach that advances the direction of arrival estimation within continual learning. CL-SSL synergizes task-specific network instantiation with progressive learning principles, addressing the diversity in microphone configurations with a novel scaling mechanism and lateral connections for knowledge retention. Demonstrated across simulated acoustic settings, CL-SSL consistently achieved high accuracy, underscoring its capacity for real-world applicability in dynamic auditory scenes. The model’s efficiency and adaptability suggest its potential as a foundational framework for future research in sound localization and acoustic signal processing, encouraging the development of sophisticated, scalable audio analysis technologies.